the sum of the squares of two positive integers is 272. The difference of the two integers is 12. Find the two integers

Answer by checkley79(3341)

X-Y=12 OR X=Y+12
(Y+12)^2+Y^2=272
Y^2+24Y+144+Y^2=272
2Y^2+24Y+144-272=0
2Y^2+24Y-128=0
2(Y^2+12-64)=0
2(Y+16)(Y-4)=0
Y-4=0
Y=4 ANS.
X=4+12X=16 ANS.
PROOF:
4^2+16^2=272
16+256=272
272=272